The surface temperature of Proxima Centauri is about 3000 K. Use Wien's law to determine λmax...

  1. The surface temperature of Proxima Centauri is about 3000 K. Use Wien's law to determine λmax in nanometers (nm). What spectral range does this wavelength correspond to (radio, visible, infrared, etc.)
Answer #1

By wein's displacement law we have

λmax T = 2898 *10^-6 m

where T = temperature in kelvin

λmax * 3000 = 2898*10^-6

λmax = 966*10^-9 m

= 966 nm

it falls in the range of ultraviolet rays